# Lexical categories of stem-forming roots in Mapudüngun verb forms

### v2

@@ 1.1.3 A Lexical Study
E.03.c ((98) b., p.: 176 in the source) is employed to illustrate a case of metonymy, yet it transpires to be a case of semantic calque from the Spanish ’las vergüenzas’. In point of fact, the online DLE dictionary provides the eighth definition for the entry "vergüenza" as follows:
@@
  1. f. pl. Órganos sexuales externos del ser humano. La pintura muestra a Adán con una hoja de parra cubriendo sus vergüenzas. (REAL ACADEMIA ESPAÑOLA: Diccionario de la lengua española, 23.ª ed., [versión 23.8 en línea]. <https://dle.rae.es> [2024-12-21].)
@@ 2 Verbal roots
As mentioned in the Introduction1, p. 1), for a computational system like the one developed here, correctly identifying the valency of verbal roots is essential for computational morphological analysis. Many of the rules governing the combination of roots and suffixes depend on this classification or on changes in value during the word formation process. This is particularly important for achieving an unambiguous analysis or minimising ambiguity [^17] In a computational morphological analysis system, ambiguity is related to the number of analyses generated by the machine. In other words, the more possible analyses the analyser returns for a lexical form, the more ambiguous the form is. Typically, the sentence context resolves the ambiguity, but this task must be handled by another system: a grammatical disambiguator. However, this tool is only helpful when analysing sentences; for isolated words, or words without context, it is less useful.
. As stated in the introduction, before undertaking the task of valency classification for Mapudüngun verbal roots, a (re)categorisation of the roots that form verbal themes in this language will be carried out. The aim is to more accurately determine the lexical category of these roots, which will help in identifying the roots to be included in the subsequent work on recognising verbal valency (Chandía, 2025). From a computational perspective, this will also enable a better morphological analysis of Mapuche verbs.
@@ 2.1 Lexical Category of Roots
The verbal roots listed by Smeets in her dictionary were examined to verify their lexical classification and to modify it if they actually belong to other lexical categories and have been verbalised instead. To achieve this, we primarily referred to the dictionaries by de Augusta, Febrés, and de Valdivia through the CORLEXIM (Chandía, 2014) platform, as well as to the autobiography of the longko [^18] Literally, longko means ‘head’, but similar to other languages, such as Catalan, it is used to denote a chief or leader. In the Mapuche context, a longko is a leader or chief in a broad sense, essentially a community guide who ensures the well-being of his lof (community). On the other hand, a toki is the chief in matters of war. In a more everyday sense, toki means ‘axe’.
Paskwal Koña [^19] The longko Paskwal Koña narrated his life to the Capuchin missionary Ernst Wilhelm von Mösbach, who transcribed it in Mapudüngun. This took place in the 1920s when Koña was still an energetic longko of 80 years. This autobiography is a unique document, not only for the information it provides about the Mapuche language and customs, but also because it is one of the few written sources in Mapudüngun containing words dictated by a Mapuche speaker.
@@
(Wilhelm et al., 1930). Examples of the roots to be verified were sought in all these sources. If any root was found used in isolation —i.e., not as part of a verbal form, nor with affixes attached— it was inferred that the original form is nominal, adjectival, or adverbial, depending on the context and usage. Many of these roots also appear in Smeets’ dictionary with the category listed in the column ‘Final Cat.’ in Table LABEL:tab1. These are non-verbal lexical categories, indicating that these words have been verbalised. In other words, the author records the meaning of the form in its original category as well as in its verbalised version. This column also reflects the lexical category reached after verifying each root in this study. The categories of the roots extracted from Smeets’ dictionary are listed in the ‘Initial Cat.’ column. Some of these roots are recorded by Smeets with an incorporated suffix, which can alter their meaning and valency when the suffix is absent. These suffixes are extracted from the original form and listed in the ‘Extracted suffixes’ column –the following Table LABEL:tab1 is only a sample available for easy verification of what is stated here. The complete table, which contains two hundred and thirty-eight roots, can be found in the Appendix, §4, Table LABEL:tab3, p. LABEL:tab3–.
@@
Table 1: Verbal Roots Reclassified As Other PoS (Sample)
Root Extracted
Suffixes
Initial
Cat.
Final
Cat.
Definition
ad- VI N the visible or characteristic aspects of something or someone, or of a group
chadi- VI N salt
dakel- VT N commission, agreement
[…]
fa- VI Dem this, here
ina- VI Av close
kachu- VI N grass, herb
la- VI Aj dead, deceased
llag- VT Aj half
machi- VI N shaman
nag- VI Av down, below, under, beneath
ngeñi- -ka- VI Aj urgent, pressing
ñochi- VI Aj peaceful, calm
paf- VI N boil, abscess
raküm- VT N closure, wall, protection
tonon- VT N weft
traf- VI Aj together, gathered, joined, fitted, interlocked
[…]
üna- VI N itch
wachi- VT N trap
yafü- VI Aj hard, firm
@@ 2.2 Valency Of Verbal Roots
The second step in the classification of roots that participate in Mapuche verbal themes is to specify the valency of verbal roots; as previously mentioned, this task will be presented in Chandía (2025). However, we will preview the results of this study and present a table of some verbal roots that co-occur with the causative suffix -(ü/e)l- (see the list of roots in p. 2.3). We also provide a list of labile roots incorporated into Düngupeyüm and a table of roots that undergo causativisation with the causative suffix -(ü)m- (Table LABEL:tab4).
2.3 Verbal Roots That Co-occur With The Causative Suffix -(ü/e)l-
In the autobiography of the longko (Wilhelm et al., 1930), twenty-one verbal roots [^20] Approximately one hundred and ten verbal roots that were not forming compounds were checked. Since this task took too long, it was decided to postpone the verification of roots when they participate in verbal compounds and exhibit a suffix in one of the forms that the causative can take: -l-, -ül-, -el-.
were found that combine with the causative -l-, as shown in the list bellow. Almost all the roots are intransitive, except for two that form both transitive and intransitive verbs, meaning they are labile or bivalent roots (for a detailed explanation see §3.4 and §3.4.1 in Chandía, 2025). In the following list, the two senses of the labile verbs are marked according to their valency, with -IV (intransitive verb) and -TV (transitive verb).
@@ 2.3 Verbal Roots That Co-occur With The Causative Suffix -(ü/e)l-
    monge- ‘-IV to live, to heal, to get healed’ / ‘-TV to revive, to heal’
@@
[…]
    ‘to wake up’
@@
    yewe- ‘-IV to feel embarrassed, to be ashamed’ / ‘-TV to respect’
@@ 2.4 Collecting Labile Roots In Düngupeyüm
In the Düngupeyüm morphological analysis system, labile roots will now appear in both the transitive and intransitive lists, which inevitably leads to increase ambiguity when analysing verb forms; however, excluding these roots would omit potential analyses. Additionally, the system has been updated to ensure that causatives, whether -(ü)m- or -(ü)l- (Chandía, 2008, §3.3 and §3.4 respectively), do not co-occur with transitive roots. Thus, causatives will only combine with labile roots in their intransitive sense. The following list comprises the labile verb roots that have been incorporated into the system:
@@
  • aye- ‘to laugh’ (-IV)
  • aye- ‘to laugh at; to laugh with’ (-TV)
  • kewa- ‘to fight, to quarrel, to struggle, to combat’ (-IV)
  • kewa- ‘to hit, to strike, to defeat, to overcome’ (-TV)
  • llüka- ‘to get frightened, to be scared, to fear, to be afraid’ (-IV)
  • llüka- ‘to scare, to frighten’ (-TV)
  • meke- ‘to take care of, to deal with, to be busy, to last, to endure’ (-IV)
  • meke- ‘to make’ (-TV)
  • monge- ‘to get life, to heal, to recover’ (-IV)
  • monge- ‘to revive, to bring back to life, to heal’ (-TV)
  • nge- ‘to be’ (-IV)
  • nge- ‘to have’ (-TV) (see nge- -IV ‘to be’, -TV ‘to have’ in Chandía, 2025)
  • püna- ‘to stick, to glue, to adhere’ (-IV)
  • püna- ‘to glue, to paste’ (-TV)
  • waychüf- ‘to fall, to get turned over, to roll, to spin’ (-IV)
  • waychüf- ‘to discard, to flip, to roll, to spin’ (-TV)
  • yewe- ‘to feel embarrassed, to be ashamed’ (-IV)
  • yewe- ‘to respect’ (-TV)
2.5 Verbal Roots That Co-occur With The Causative Suffix -(ü)m-
In Chandía, 2025: §3.3, it will be shown that verifying the presence of the causative suffix -(ü)m- in Mapuche verb forms is a reliable method for determining that the root carrying it is intransitive. However, it is also necessary to verify whether the root is genuinely verbal or it belongs to another lexical category, a task that this document addresses. Furthermore, as noted by Longkon (2011a), Golluscio (2007) and Smeets, not all intransitive roots accept causativisation with -(ü)m-. Therefore, it is important to examine these authors’ claims regarding the role of telicity or atelicity in the selection of the causative suffix. In Table LABEL:tab4, all the roots that have been verified to combine with the causative suffix -(ü)m- are presented, the following Table LABEL:tab2 is only a sample available for easy verification of what is stated in Chandía, 2008: §3.3.
Table 2: Roots That Causativise With -(ü)m- (Sample)
@@ 2.5 Verbal Roots That Co-occur With The Causative Suffix -(ü)m-
ad N form, place, familiar, face, image, custom ad-üm- to be experienced/skilled (in something) K. S.
anü VI to get settled, to get sit anü-m- to settle, to set, to put, to place K. S. G.
fa Dem this, here fa-m- to be like this K. S.
firkü Aj fresh, cool firkü-m- to cool, to warm slightly, to take the chill off K.
kon VI to enter, to begin kon-üm- to make come in, to add K. G.
la Aj dead, deceased lang-üm- to kill K. S. G.
(l)el VI to let oneself, to become loose, to detach (l)el-üm- to release, to let go, to kill K.
llum VI to get hidden, to conceal oneself llum-üm- to hide, to conceal K.
[…]
nor Aj straight, correct, right nor-üm- to correct, to rectify, to straighten S.
pewü VI to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m- to measure, to reach out using the arms, to twist, to turn K.
pümü Aj tense, tight, taut, strained pümü-m- to stretch, to tighten, to tense K.
puw VI to arrive, to reach, to achieve, to stay, to remain puw-üm- to send, to deliver, to convey, to serve K.
trar VI to ooze, to suppurate, to secrete trar-üm- to drain the pus S.
tüng VI to calm down, to take time tüng-üm- to leave alone, not to disturb K.
üy N name üy-üm- to name, to denominate S.
wim Aj accustomed, used to, meek, tame, domesticated wim-üm- to make accustomed, to domesticate, to tame S.
yung Aj pointed, sharp yung-üm- to sharpen, to make pointed K.
@@ 3 Conclusions
In the present work, we undertook a reclassification of the lexical categories of roots involved in verbal themes of the Mapuche language, which are part of the inventory of examples in the descriptive grammar of Mapudüngun developed by the anthropologist Ineke Smeets. This is a preliminary task to valency classification (Chandía, 2025), the study that motivated us to begin this lexical category review. The task of valency classification of Mapuche verbal roots required an accurate classification of the lecixal categories of the roots to avoid attributing verbal features, such as valency, to elements that cannot exhibit them because they are not verbal. A simple verification method for the lexical category was devised, as explained in §2.1, p.2.1, and despite its simplicity, it has proven to be effective, with results shown in Table LABEL:tab3, p. LABEL:tab3).
@@
In a computational system like an automatic morphological analyser, where it is crucial to have specific and non-redundant data, ambiguity can be significantly reduced if roots are listed, as much as possible, under a single classification per form. Despite this, we chose to list roots with ambivalent behaviour (see §2.4, p. 2.4) in both lists, intransitive verbal roots and transitive verbal roots, with each list reflecting the root’s valency-related sense. This approach might introduce varying degrees of ambiguity in the analyses produced by the machine. However, besides the ambiguity, it also allows for potential analyses that would be excluded if ambivalent roots were not collected in this manner.
It is essential to first identify the category most likely to be the source of the multiple interpretations or derivations that a form, used as the main component of the verbal theme, might generate (see §2.1, p. 2.1). Therefore, the primary task is to ensure that valency values are sought in inherently verbal roots rather than in verbalised roots derived from other categories (see Table LABEL:tab3, p. LABEL:tab3). This can be challenging in a language like Mapudüngun, as there is often insufficient material or stored knowledge to determine the etymology of much of its vocabulary, largely due to historical and social factors. Techniques like the one employed in this study, while reflecting a highly conditioned perspective, often diverge from fully verifiable data. Nonetheless, it is at least known that Wilhelm et al. (1930) is one of the best references for the Mapuche language (see footnote 19, p. 19), which was the main reason for its use in this research.
To conclude, the decision to publish the study in two parts —lexical classification presented here and valency classification presented in Chandía, 2025— aims to clarify the findings. As stated in the summary, ‘‘The outcomes of this categorical and valency review directly benefit the computational analyser, as they are implemented as soon as they are verified. Additionally, these findings are expected to enhance the understanding of the Mapuche language’’. Readers are invited to review the accompanying article to this first work, which provides a comprehensive understanding of the results of this research.

### v3

@@ 1.1.3   A Lexical Study
E.03c ((98) b., p.: 176 in the source) is employed to illustrate a case of metonymy, yet it transpires to be a case of semantic calque from the Spanish ‘las vergüenzas’. In point of fact, the online DLE dictionary provides the eighth definition for the entry “vergüenza” as follows:
@@
f. pl. Órganos sexuales externos del ser humano. La pintura muestra a Adán con una hoja de parra cubriendo sus vergüenzas. (REAL ACADEMIA ESPAÑOLA: Diccionario de la lengua española, 23.ª ed., [versión 23.8 en línea]. <https://dle.rae.es> [2024-12-21].)
@@
1.2   Summary of the Work
This study is a systematic reclassification of the lexical categories of roots involved in Mapudüngun verbal themes. The primary goal of the work is to establish a reliable foundation for valency classification (Chandía, 2026) by ensuring that roots identified as verbal are indeed inherently verbal, rather than verbalised forms belonging to other lexical categories (e.g., nouns, adjectives, adverbs).
The methodology employed is a straightforward verification process: roots listed as verbal in Smeets (2008)’s dictionary are cross-referenced against their usage in isolation in historical dictionaries (de Augusta, Febrés, de Valdivia) and the autobiography of the longko Paskwal Koña (Wilhelm et al., 1930). If a root is found used without derivational suffixes in these sources, it is inferred that its original category is non-verbal. This process yielded a comprehensive table (Table 3 in the 4   Appendix) showing 238 roots initially classified as verbs but reclassified as nouns, adjectives, adverbs, or demonstratives.
Key findings from this study include:
  • A significant number of roots in Smeets (2008)’s verbal inventory are verbalised forms of other lexical categories.
  • Lists of verbal roots that co-occur with the causative suffix -(ü/e)l and -(ü)m were compiled from the longko Koña’s memoirs (Wilhelm et al., 1930).
  • A set of labile roots (those that can be used both transitively and intransitively without morphological change) was identified.
The following article (Chandía, 2026) builds directly on this foundation. The lexical classifications established here are taken as the starting point for the analysis that follows.
@@ 2   Verbal Roots
As mentioned in the 1   Introduction1), for a computational system like the one developed here, correctly identifying the valency of verbal roots is essential for computational morphological analysis. Many of the rules governing the combination of roots and suffixes depend on this classification or on changes in value during the word formation process. This is particularly important for achieving an unambiguous analysis or minimising ambiguity [^14] In a computational morphological analysis system, ambiguity is related to the number of analyses generated by the machine. In other words, the more possible analyses the analyser returns for a lexical form, the more ambiguous the form is. Typically, the sentence context resolves the ambiguity, but this task must be handled by another system: a grammatical disambiguator. However, this tool is only helpful when analysing sentences; for isolated words, or words without context, it is less useful.. As stated in the introduction, before undertaking the task of valency classification for Mapudüngun verbal roots, a (re)categorisation of the roots that form verbal themes in this language will be carried out. The aim is to more accurately determine the lexical category of these roots, which will help in identifying the roots to be included in the subsequent work on recognising verbal valency (Chandía, 2026). From a computational perspective, this will also enable a better morphological analysis of Mapuche verbs.
@@ 2.1   Lexical Category of Roots
The verbal roots listed by Smeets (2008) in her dictionary were examined to verify their lexical classification and to modify it if they actually belong to other lexical categories and have been verbalised instead. To achieve this, we primarily referred to the dictionaries by de Augusta, Febrés, and de Valdivia through the CORLEXIM (Chandía, 2014) platform, as well as to the autobiography of the longko [^15] Literally, longko means ‘head’, but similar to other languages, such as Catalan, it is used to denote a chief or leader. In the Mapuche context, a longko is a leader or chief in a broad sense, essentially a community guide who ensures the well-being of his lof (community). On the other hand, a toki is the chief in matters of war. In a more everyday sense, toki means ‘axe’. Paskwal Koña [^16] The longko Paskwal Koña narrated his life to the Capuchin missionary Alois Wilhelm von Mösbach (padre Ernesto), who transcribed it in Mapudüngun. This took place in the 1920s when Koña was still an energetic longko of 80 years. This autobiography is a unique document, not only for the information it provides about the Mapuche language and customs, but also because it is one of the few written sources in Mapudüngun containing words dictated by a Mapuche speaker. (Wilhelm et al., 1930). Examples of the roots to be verified were sought in all these sources. If any root was found used in isolation —i.e., not as part of a verbal form, nor with affixes attached— it was inferred that the original form is nominal, adjectival, or adverbial, depending on the context and usage. Many of these roots also appear in Smeets (2008)’s dictionary with the category listed in the column ‘Final Cat.’ in Table 1. These are non-verbal lexical categories, indicating that these words have been verbalised. In other words, the author records the meaning of the form in its original category as well as in its verbalised version. This column also reflects the lexical category reached after verifying each root in this study. The categories of the roots extracted from Smeets (2008)’s dictionary are listed in the ‘Initial Cat.’ column. Some of these roots are recorded by Smeets (2008) with an incorporated suffix, which can alter their meaning and valency when the suffix is absent. These suffixes are extracted from the original form and listed in the ‘Extracted suffixes’ column —the following Table 1 is only a sample available for easy verification of what is stated here. The complete table, which contains two hundred and thirty-eight roots, can be found in the 4   Appendix, §4, Table 3.
@@
?tablename? 1: Verbal Roots Reclassified as Other Parts of Speech (Sample)
Root Extracted Initial Final Definition
Suffixes Cat. Cat.
ad- IV N the visible or characteristic aspects of something or someone, or of a group
chadi- IV N salt
dakel- TV N commission, agreement
echid- $\sim$ echiw- IV N sneeze
fa- IV Dem this, here
ina- IV Av close
kachu- IV N grass, herb
la- IV Aj dead, deceased
llag- TV Aj half
machi- IV N shaman
nag- IV Av down, below, under, beneath
ngeñi- -ka IV Aj urgent, pressing
ñochi- IV Aj peaceful, calm
paf- IV N boil, abscess
raküm- TV N closure, wall, protection
tonon- TV N weft
traf- IV Aj together, gathered, joined, fitted, interlocked
uma- $\sim$ umañ- $\sim$ umag- $\sim$ umaw- IV N sleepiness, drowsiness
üna- IV N itch
wachi- TV N trap
yafü- IV Aj hard, firm
Abbreviations: IV = Intransitive Verb; TV = Transitive Verb; N = Noun; Aj = Adjective; Av = Adverb;
Dem = Demonstrative.
@@ 2.2   Valency of Verbal Roots
The second step in the classification of roots that participate in Mapuche verbal themes is to specify the valency of verbal roots; as previously mentioned, this task will be presented in Chandía (2026). However, we will preview the results of this study and present a list of some verbal roots that co-occur with the causative suffix -(ü/e)l (see §2.3). We also provide a list of labile roots (see §2.4) and a sample table of roots that undergo causativisation with the causative suffix -(ü)m (Table 2).
2.3   Verbal Roots That Co-occur With The Causative Suffix -(ü/e)l
In the autobiography of the longko (Wilhelm et al., 1930), twenty-one verbal roots [^17] Approximately one hundred and ten verbal roots that were not forming compounds were checked. Since this task took too long, it was decided to postpone the verification of roots when they participate in verbal compounds and exhibit a suffix in one of the forms that the causative can take: -l, -ül, -el. were found that combine with the causative -l, as shown in the list below. Almost all the roots are intransitive, except for two that form both transitive and intransitive verbs, meaning they are labile or bivalent roots (for a detailed explanation see the companion valency article, Chandía, 2026). In the following list, the two senses of the labile verbs are marked according to their valency, with IV (intransitive verb) and TV (transitive verb).
@@ 2.3   Verbal Roots That Co-occur With The Causative Suffix -(ü/e)l
    monge- IV ‘to live, to heal, to get healed’ / TV ‘to revive, to heal’
@@
    nepe- [^18] The computational system generates all possible analyses, which sometimes leads to conjectures about the origin of certain Mapudüngun forms. For instance, nepe- meaning ‘to wake up’ might have originated from the compound ne- $\sim$ nie- + pe: ‘to have’ + ‘sight, vision’. Similarly, it can be inferred that pewma meaning ‘dream’ comes from pe- + uma(g): ‘sight, vision’ + ‘sleep’. ‘to wake up’
@@
    yewe- IV ‘to feel embarrassed, to be ashamed’ / TV ‘to respect’
@@ 2.4   Collecting Labile Roots In Düngupeyüm
In the Düngupeyüm morphological analysis system, labile roots will now appear in both the transitive and intransitive lists, which inevitably leads to increased ambiguity when analysing verb forms; however, excluding these roots would omit potential analyses. Additionally, the system has been updated to ensure that causatives, whether -(ü)m or -(ü)l, do not co-occur with transitive roots. Thus, causatives will only combine with labile roots in their intransitive sense. The following list comprises the labile verb roots that have been incorporated into the system:
@@
  1. 1.
    aye- ‘to laugh’ (IV)
    aye- ‘to laugh at; to laugh with’ (TV)
  2. 2.
    kewa- ‘to fight, to quarrel, to struggle, to combat’ (IV)
    kewa- ‘to hit, to strike, to defeat, to overcome’ (TV)
  3. 3.
    llüka- ‘to get frightened, to be scared, to fear, to be afraid’ (IV)
    llüka- ‘to scare, to frighten’ (TV)
  4. 4.
    meke- ‘to take care of, to deal with, to be busy, to last, to endure’ (IV)
    meke- ‘to make’ (TV)
  5. 5.
    monge- ‘to get life, to heal, to recover’ (IV)
    monge- ‘to revive, to bring back to life, to heal’ (TV)
  6. 6.
    nge- ‘to be’ (IV)
    nge- ‘to have’ (TV) (see §3.4.2.3 in Chandía, 2026)
  7. 7.
    püna- ‘to stick, to glue, to adhere’ (IV)
    püna- ‘to glue, to paste’ (TV)
  8. 8.
    waychüf- ‘to fall, to get turned over, to roll, to spin’ (IV)
    waychüf- ‘to discard, to flip, to roll, to spin’ (TV)
  9. 9.
    yewe- ‘to feel embarrassed, to be ashamed’ (IV)
    yewe- ‘to respect’ (TV)
2.5   Verbal Roots That Co-occur With The Causative Suffix -(ü)m
In the companion valency article (Chandía, 2026): §3.3), it will be shown that verifying the presence of the causative suffix -(ü)m in Mapuche verb forms is a reliable method for determining that the root carrying it is intransitive. However, it is also necessary to verify whether the root is genuinely verbal or belongs to another lexical category, a task that this document addresses. Furthermore, as noted by Longkon (2011), Golluscio (2007) and Smeets (2008), not all intransitive roots accept causativisation with -(ü)m. Therefore, it is important to examine these authors’ claims regarding the role of telicity or atelicity in the selection of the causative suffix. In Table 4 (4   Appendix), all the roots that have been verified to combine with the causative suffix -(ü)m are presented. The following Table 2 is only a sample available for easy verification of what is stated in the valency article.
?tablename? 2: Roots That Causativise With -(ü)m (Sample)
@@ 2.5   Verbal Roots That Co-occur With The Causative Suffix -(ü)m
ad N form, place, familiar, face, image, custom ad-üm to be experienced/skilled (in something) K. S.
anü IV to get settled, to get sit anü-m to settle, to set, to put, to place K. S. G.
fa Dem this, here fa-m to be like this K. S.
firkü Aj fresh, cool firkü-m to cool, to warm slightly, to take the chill off K.
kon IV to enter, to begin kon-üm to make come in, to add K. G.
la Aj dead, deceased lang-üm to kill K. S. G.
(l)el IV to let oneself, to become loose, to detach (l)el-üm to release, to let go, to kill K.
llum IV to get hidden, to conceal oneself llum-üm to hide, to conceal K.
nel Aj loose, untied, unleashed nel-üm $\sim$ nelk-üm to release, to let go, to free S.
nor Aj straight, correct, right nor-üm to correct, to rectify, to straighten S.
pewü IV to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m to measure, to reach out using the arms, to twist, to turn K.
pümü Aj tense, tight, taut, strained pümü-m to stretch, to tighten, to tense K.
puw IV to arrive, to reach, to achieve, to stay, to remain puw-üm to send, to deliver, to convey, to serve K.
trar IV to ooze, to suppurate, to secrete trar-üm to drain the pus S.
tüng IV to calm down, to take time tüng-üm to leave alone, not to disturb K.
üy N name üy-üm to name, to denominate S.
wim Aj accustomed, used to, meek, tame, domesticated wim-üm to make accustomed, to domesticate, to tame S.
yung Aj pointed, sharp yung-üm to sharpen, to make pointed K.
aThe “Collected from” column indicates which source contains examples of the root with the causative suffix.
K. = Wilhelm et al., 1930 (autobiography of longko Paskwal Koña); S. = Smeets, 2008 (A Grammar of Mapuche); G. = Golluscio, 2007 (Gramática descriptiva del mapuche).
@@ 3   Conclusions
In the present work, we undertook a reclassification of the lexical categories of roots involved in verbal themes of the Mapuche language, which are part of the inventory of examples in the descriptive grammar of Mapudüngun developed by the anthropologist Ineke Smeets (2008). This is a preliminary task to valency classification (Chandía, 2026), the study that motivated us to begin this lexical category review. The task of valency classification of Mapuche verbal roots required an accurate classification of the lexical categories of the roots to avoid attributing verbal features, such as valency, to elements that cannot exhibit them because they are not verbal. A simple verification method for the lexical category was devised, as explained in §2.1, and despite its simplicity, it has proven to be effective, with results shown in Table 3 (4   Appendix).
@@
In a computational system like an automatic morphological analyser, where it is crucial to have specific and non-redundant data, ambiguity can be significantly reduced if roots are listed, as much as possible, under a single classification per form. Despite this, we chose to list roots with ambivalent behaviour (see §2.4) in both lists, intransitive verbal roots and transitive verbal roots, with each list reflecting the root’s valency-related sense. This approach might introduce varying degrees of ambiguity in the analyses produced by the machine. However, besides the ambiguity, it also allows for potential analyses that would be excluded if ambivalent roots were not collected in this manner.
It is essential to first identify the category most likely to be the source of the multiple interpretations or derivations that a form, used as the main component of the verbal theme, might generate (see §2.1). Therefore, the primary task is to ensure that valency values are sought in inherently verbal roots rather than in verbalised roots derived from other categories (see Table 3 in the 4   Appendix). This can be challenging in a language like Mapudüngun, as there is often insufficient material or stored knowledge to determine the etymology of much of its vocabulary, largely due to historical and social factors. Techniques like the one employed in this study, while reflecting a highly conditioned perspective, often diverge from fully verifiable data. Nonetheless, it is at least known that Wilhelm et al. (1930) is one of the best references for the Mapuche language (see footnote 16), which was the main reason for its use in this research.
To conclude, the decision to publish the study in two parts —lexical classification presented here and valency classification presented in Chandía, 2026— aims to clarify the findings. As stated in the summary, “The outcomes of this categorical and valency review directly benefit the computational analyser, as they are implemented as soon as they are verified. Additionally, these findings are expected to enhance the understanding of the Mapuche language”. Readers are invited to review the accompanying article to this first work, which provides a comprehensive understanding of the results of this research.